OpenStudy (xxcecelovxx):

What is the surface area of a sphere with a radius of 60 m in terms of π?

OpenStudy (xxcecelovxx):

https://gyazo.com/9164cf47aed8f59e14ab276218d4edd6

Vocaloid (vocaloid):

Surface area = 4pi * r^2

OpenStudy (xxcecelovxx):

45238.9342117

OpenStudy (xxcecelovxx):

?

Vocaloid (vocaloid):

leave it in terms of pi

jabez177 (jabez177):

Isn't that Volume?

Vocaloid (vocaloid):

no

OpenStudy (xxcecelovxx):

how do you get it in terms of pi?

Vocaloid (vocaloid):

volume is 4/3 pi r ^3

Vocaloid (vocaloid):

"in terms of pi" just means multiply out everything except pi, then write pi next to it

Vocaloid (vocaloid):

so

Vocaloid (vocaloid):

4pi * r^2 = 4 * pi * 60^2 = 14400 pi

Vocaloid (vocaloid):

and that's your answer

OpenStudy (xxcecelovxx):

thank you very much!